If a human body has a total surface area of 1.8 m, what is the total force on the body due to the atmosphere at sea level (1.01 x 10 Pa)?
Answer #1

We have inward pressure to balance outward pressure .

At every point of the body the atmosphere pressure is

P=1.01×10^5 Pa=1.01×10^5 N/m^2

We know that

Force =F=PA . Here A=1.8m^2

F=1.01×10^5 ×1.8

F=1.818 ×10^5 N Answer.